The bass in Clear Lake have weights that are normally distributed with a mean of 2.3 pounds and a standard deviation of 0.6 pounds. What percentage of all randomly caught groups of 3 bass should weigh between 2.1 and 2.6 pounds?

Enter your answer as a percentage rounded to one decimal place.
Answer #1

Solution :

Given that ,

mean = = 2.3

standard deviation = = 0.6

n = 3

= 2.3

=  / n= 0.6 / 3=0.345

P(2.1<     < 2.6) = P[(2.1-2.3) /0.345 < ( - ) /   < (2.6-2.3) /0.345 )]

= P( -0.58< Z <0.87 )

= P(Z < 0.87) - P(Z <-0.58 )

Using z table

=0.8078-0.2810

=0.5268

=52.7%